# KEN KANGDA WREN1

### v4

@@ 3.5.1 Momentum Effect and Status Quo Bias
Conditional excess (41) exhibits Momentum, a key price-anomaly (Jegadeesh and Titman (1993)): if $\rho K^{\pm 1}\gg 1$, excess-to-date ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction. The effect applies to $\rho K^{\pm 1}\ll 1$ also: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching here defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only under $\rho\approx 1$ (little model-uncertainty).
The effect grows with $\rho$, which by definition (40) measures bias against change. From (41), momentum trading earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm})^{\frac{1}{2}}-1}{(\rho K^{\pm})^{\frac{1}{2}}+1}S^{\Delta}.$ (43)
Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momentum $v^{\pm mo}_{max}$. As a function of bias $\rho$, risk-reward curve (41) and its peak (43) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as seen in Ang et al. (2006).
@@ 3.5.2 Low-Risk Effect and Status Quo Bias
Low-Risk Effect (e.g. Ang et al. (2006)) turns the classical dictum ’high risk high reward’ on its head. Momentum risk-reward (41) already hints at it: peak-reward (43) is in the low-volatility region if bias $\rho$ is sizeable. The effect is clearer under volatility-conditioning (42), whose peak-location and -size can be found by routine methods (Appendix C). The solution at $\rho=1$ (no bias) and leading-order solutions at $\rho\gg 1$ (high bias) are revealing and may be given by the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are:
@@
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
There is otherwise a low-risk effect: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, corresponding to low risk if $\rho\gg 1$; peak volatility, $v=\frac{1}{2}$, is rewarded poorly ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$. Note the separation of risk-pricing $K$ and bias $\rho$: the latter is revealed by peak-location $v^{\sigma}_{max}$, while the former, by peak-reward $rp^{\sigma}_{max}$, which is that set ex-ante for peak-risk (Remark 9). No external rules are needed to tell ’intended pricing’ from ’unintended bias’.
@@ 4 Summary and Discussion
We investigate the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived, exposing model-risk as a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. It offers an integrated and practicable approach to these highly relevant and interconnected topics.
We make use of a feature of standard hypothesis testing (Lemma 1) that ’informationally equivalent’ tests must be essentially identical. This makes risk-neutral equivalent formulations under model-risk path-dependent and intractable. On the other hand, it limits viable economic asset-pricing (Remark 1) under model-risk to just one form (Proposition 1), leading to an intuitive risk-pricing formula with intuitive properties (Corollary 1).
The sole risk-pricing parameter is associated with a conservation law of inference and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) model-risk pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$ (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 9).
The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$ (40), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’. It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.5.1-3.5.2).
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as models are given by probability laws that are equivalent before risk-resolution. The theoretical and practical aspects of this topic will be examined in a follow-up study.
Our approach may be extended to model-risks characterisable as $(N+1)$-tuples of independent binary outcomes ${B}^{N+1}:=(B_{0}B_{1}...B_{N})$, $N\in\mathbb{N}$, through iteration. It may be applied first to the $2^{N}$ pairs of ${B}^{N+1}$-sure, elementary, price processes, then to the resulting $2^{N-i}$ pairs of ${B}^{N+1-i}$-sure composites, $i=1,...,N$, each discounting the $i$-tuple model-risk $(B_{0}B_{1}...B_{i-1})$.
@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Any continuous process with independent increments are essentially Wiener. By definition (Item-4, Section 1.1), the log-LR process under i.i.d data, in terms of some constant $\sigma^{l}>0$ (the ’signal-to-noise’ of data), reads:
@@
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Time-variation may be modelled via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$, again Wiener and with absolutely continuous time-varying parameters (e.g. Kallsen (2006)), along with inferential beliefs, an Ito process of it ((3)), read as follows:
    $dl^{+/-}_{t}(B)=\mu^{l^{+/-}}_{t}(B)dt+\sigma^{l}_{t}dw_{t},\text{ }\text{ }\mu^{l^{+/-}}_{t}(B):=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2};$ (A.2)
@@
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In usual situations (e.g. i.i.d data), risk-resolution occurs at infinity. Clock-change allows continuous-time regular inference to have finite resolution-times, by compactifying uniform regular inference to bring resolution $T_{\tau}(B)=\infty$ forward to some $T(B):=\tau^{-1}(\infty)<\infty$. The resulting log-LR process diverges as $t\to T(B)$, but the associated beliefs and so asset-pricing remain finite and continuous under Bayes’ Rule (3).
@@
    Complete Markets and Unhedgeable Model-Risks. Market completeness excludes unpredictable resolutions (price-jumps) and so demands regular inference. In this setting, to be unhedgeable and thus priced economically (Remark 1), model-risk $B$ cannot be resolvable before horizon $T$ (so $T(B)>T$); this is the default case for model-risk studies. The mixed case of probabilistic resolution, ($\mathbb{P}\big{(}\{T(B)\leq t+s\}|F_{t}\big{)}\in(0,1)$, $\forall t<T$, $\forall s\in(0,T-t]$, which may occur in continuous-time complete markets so long as $T(B)$ is predictable, involves treating the unhedgeable part of the model-risk only.
@@
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely, and hence a finite total variance ((A.2)). That is, the non-resolution of $B$-outcomes (Item-4, Section 1.1) equates to (A.2-A.3) satisfying Novikov’s Condition, and vice versa.
  5. 5.
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to T(B)$, since:
@@
    where the RHS is convergent under equivalence.
@@
    No adjacent log-LR processes, unless identical up to a constant difference, can be informationally equivalent, that is, be deterministically and time-homogenously related to one another, nor can any inferential processes derived from them.
    Proof. It is a feature of inference (and due to Bayes’ Rule). It can be verified concretely via Ito’s Lemma here, given (A.2) and restrictions to twice-differentiable maps between processes (as in Proposition 1). Any such mapping $g:{l}^{+/-}_{t}\in\mathbb{R}\mapsto\hat{l}^{+/-}_{t}=g(l^{+/-}_{t})\in\mathbb{R}$, $t\in(0,T(B))$, must solve the ordinary differential equation:
@@
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$ and so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; divergence $\lim_{t\to T(+)}\hat{l}^{+/-}_{t}=\infty$ requires $g^{\prime}(0)=1$; If $B=-$, then $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$ and $\lim_{t\to T(-)}\hat{l}^{+/-}_{t}=-\infty$ imply the same.
    Remark 14.
    Given identical LR up to a constant factor, inference process $\{\hat{\pi}^{+}_{t}\}$, a $(\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B})$-martingale by definition, is also a $(\hat{\pi}^{B}_{0}\mathbf{W}_{B})$-martingale ((3)), meaning $\hat{\mathbf{W}}_{B}=\mathbf{W}_{B}$ (Martingale Representation Theorem and so completeness of the Wiener basis): thus, only tests based on identical underlying beliefs can be informationally equivalent.
@@ B PROOF OF PROPOSITION 1
Proof. Part I. Viable model-risk only pricing $\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}_{t}\}$, given twice-differentiable coefficients $\{A^{+}_{t}(\pi^{+}_{t})\}$, must have the canonical form.
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d(\mathbf{W}^{T}_{-})_{t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-2&4, Section 1.1), we have:
$dS^{(0|A)}_{t}=\big{[}($ $\mathbf{1}_{+}-A^{+}_{t})r^{Z\Delta}_{t}dt\text{ }+\text{ }\sigma^{Z}_{t}dw^{Z}_{t}\big{]}\text{ }+\text{ }$ (B.1)
$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw^{\prime}_{t}\bigg{]},$ (B.2)
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’ ratio) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), and $\sigma^{lD}_{t}$ likewise, from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw^{\prime}_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
@@
$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (B.3)
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$: we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$ and the following relationship, by (8-11) and (19),
@@
Reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and any RNE $\{\hat{\pi}^{B}_{t}\}$ are ’adjacent’ (Item-5, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23-24) when $\{{\hat{\pi}}^{B}_{t}\}=\{\Pi^{B}_{t}\}$, we have $\mathcal{O}({\pi}^{B}_{t}-{\hat{\pi}}^{B}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk ((11)). In turn, by Lemma 1, to balance the LHS and RHS of (B.4), we have: $\mathcal{O}({\hat{RP}^{B}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}(\partial_{\pi}A^{B}_{t})=\mathcal{O}(\partial^{2}_{\pi}A^{B}_{t})=\mathcal{O}(1)$ and $\mathcal{O}(d({\hat{RP}^{B}_{t})})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
@@
Consider first the case of trivial $\{D_{t}\}$. The drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$, where $d(\hat{W}^{Z}_{B})_{t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d(\hat{RP}^{B}_{t})$, $t<T$.
Being the RNE measure, the $(\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T})$-expectation of (B.1) must offset that of (B.2). Write that of (B.1) as $\nu_{t}:=(\hat{\pi}^{+}_{t}-A^{+}_{t})r^{Z\Delta}_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}\hat{r}_{B,\text{ }t}$ and note $\hat{\pi}^{+}_{t}-\pi^{+}_{t}=(\hat{\pi}^{+}_{t}-A^{+}_{t})-(\pi^{+}_{t}-A^{+}_{t})$, viability thus requires:
$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}-({\pi}^{+}_{t}-A^{+}_{t})-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
@@
By Remark 15, we have $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$ and $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side of the above vanishes on its own, meaning:
$\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)

### v6

@@ 3.4.1 Momentum Effect and Status Quo Bias
Conditional excess (40) exhibits Momentum, the well-known anomaly (Jegadeesh and Titman (1993)): for $\rho\gg 1$, any excess-to-date $\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction. The effect grows with $\rho$, which by its definition (38) is an indicator of bias against change.
Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only when $\rho=1$ (e.g. when there is no model-uncertainty).
Momentum trading (40) earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm 1})^{\frac{1}{2}}-1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1}S^{\Delta}.$ (42)
Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momenta $v^{\pm mo}_{max}$. As a function of bias $\rho$, risk-reward (40) and its peak (42) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as observed in Ang et al. (2006).
@@ 3.4.2 Low-Risk Effect and Status Quo Bias
Low-Risk Effect (see e.g. Ang et al. (2006)) turns the classical dictum ’high risk high reward’ on its head. Momentum risk-reward (40) already hints at it: peak-reward (42) is in the low-volatility zone for large bias $\rho\gg 1$. The effect is stronger under volatility-conditioning (41), whose peak-location and -size can be found by usual methods (Appendix C). The solution at $\rho=1$ (no bias) and the leading-order solutions at $\rho\gg 1$ (high bias) are revealing and happen to have the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are given by:
@@
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias. There is a low-risk effect otherwise: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, while peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
Note the separation: bias $\rho$ is revealed by peak-location $v^{\sigma}_{max}$, and risk-pricing $K$, by peak-reward $rp^{\sigma}_{max}$, which is exactly that set ex-ante for peak-risk (Remark 7). No external speculation is needed to tell ’intended pricing’ from ’unintended bias’.
@@ 4 Summary and Discussion
We study the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
The results show that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. They offer an integrated and practicable approach to these highly relevant and interconnected topics.
We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1).
The sole risk-pricing parameter is associated with a conserved quantity of inferential dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-to-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 7).
The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$, (38), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’. It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.4.1-3.4.2).
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution (that is predictable). The theoretical and practical aspects of this topic will be examined in a follow-up study.
@@ A Properties of Inferential Hypothesis Testing
    Standard Continuous-Time Inference. Given the setup and Item-5 of Section 1.1, the log-LR process of time-homogeneous Wiener data processes has the form below (Peskir and Shirayev (2006)), in terms of some constant $\sigma^{l}>0$ characterising the signal-to-noise ratio of the given data,
@@
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Any time-variation may be introduced via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is again Wiener, but with continuous time-varying parameters (e.g. Kallsen (2006)). It and its resulting inferential belief process under Bayes’ Rule (3) read: in terms of some standard Wiener process $\{w_{t}\}$,
    $dl^{+/-}_{t}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2}+\sigma^{l}_{t}dw_{t};$ (A.2)
@@
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In standard situations, with i.i.d data, risk-resolution occurs only at infinity. Predictable finite-time resolution can be modelled via a clock-change that brings resolution forward to some $T_{\mathcal{B}}\leq\infty$. The resulting log-LR process diverges as $t\to T_{\mathcal{B}}$, but the associated beliefs and so asset-pricing remain finite and continuous at $T_{\mathcal{B}}$ under Bayes’ Rule (3).
@@
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely and hence a finite total variance for the associated log-LR process of form (A.2). The resolution of $B$-outcomes (Item-5, Section 1.1) thus equates to the associated LR process (the exponentiation of (A.2)), or equivalently, (A.3), violating Novikov’s Condition, and vice versa.
@@
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to\infty$, since:
@@
    where the RHS is convergent under equivalence; this is the case even when each of $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ is resolving (i.e. divergent in total variance and so total drift).
@@
    No two resolving log-LR processes, unless identical up to a constant difference, can be deterministically and time-homogenously related to one another.
[…]
@@
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$, so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; we then have $g^{\prime}(0)=1$ as $\{\hat{l}^{+/-}_{t}\}$ and $\{{l}^{+/-}_{t}\}$ are resolving. If $B=$-, so $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$, then almost-sure divergence $\lim_{t\to\infty}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to\infty}{l}^{+/-}_{t}$ requires the same.
@@ B PROOF OF PROPOSITION 1
[…]
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-3&5, Section 1.1), we have:
[…]
$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}\cdot(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw_{t}\bigg{]},$ (B.2)
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-2, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
@@
$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (B.3)
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
@@
Reference model-risk beliefs $\{{\pi}^{+}_{t}\}$ and any of its RNE $\{\hat{\pi}^{+}_{t}\}$ are adjacent (Item-4, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23) when $\{{\hat{\pi}}^{+}_{t}\}=\{\Pi^{+}_{t}\}$, we have $\mathcal{O}({\pi}^{+}_{t}-{\hat{\pi}}^{+}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk (11). In turn, by Lemma 1, the LHS and RHS of (B.4) being equal implies $\mathcal{O}({\hat{RP}^{+}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}((A^{+}_{t})^{\prime}_{\pi})=\mathcal{O}((A^{+}_{t})^{\prime\prime}_{\pi})=\mathcal{O}(1)$ and $\mathcal{O}(d{\hat{RP}^{+}_{t}})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
@@
Consider first the case of trivial $\{D_{t}\}$, so that the drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$ satisfying $d\hat{W}^{Z}_{B,t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d\hat{RP}^{B}_{t}$, $t<T$.
[…]
$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(A^{+}_{t}-{\pi}^{+}_{t})+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
@@
By Remark 12, we have $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side vanishes on its own, meaning:
$\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)